# Online Makespan Scheduling under Scenarios

### v1

@@ Abstract
For varying values of the number of scenarios and machines, we explore the competitiveness of online algorithms. We prove tight and near-tight bounds, several of which are achieved through novel constructions. In particular, we leverage the interplay between the unit processing time case of our problem and the hypergraph coloring problem both ways: We use hypergraph coloring techniques to steer an adversarial family of instances proving lower bounds for our problem, which in turn leads to lower bounds for several variants of online hypergraph coloring.
@@ 1 Introduction
Moreover, even special cases of online makespan scheduling under scenarios correspond to interesting questions in their own regard. The special case $p_{j}\equiv 1$ of unit processing times is related to some cases of discrepancy minimization as well as hypergraph coloring.
In the large array of previous research on combinatorial optimization under scenarios, it was shown for several offline problems that optimizing under a number of scenarios is computationally harder than their underlying single-scenario counterparts. Accordingly, our goal in this work is to gain a better understanding of the intersection of scenario-based models and competitiveness, in particular how competitiveness diminishes under a growing number of scenarios.
@@ 1.1 Our Contribution
By a simple example, one can show that for any number $m\geq 2$ of machines and (at least) $3$ scenarios, no algorithm with a competitive ratio better than $2$ can exist, even for jobs with unit processing times. To underline the competitiveness gap between two and three scenarios, we contrast the aforementioned result with an algorithm for two machines and two scenarios, which is one of our main results.
@@
This algorithm as well as its analysis combine techniques well-known from algorithms for online makespan scheduling with novel ideas that concern the scenario-based properties of a schedule. We first alter the definition of competitive ratio by allowing only specific lower bounds for an offlie optimum, which allows us to make structural modifications and restrictions on worst-case instances. What sets our analysis apart from previous work is that we alternate between forward and reverse engineering: Assuming that our algorithm obeys some fixed rules, we explore what we may assume about a well-structured worst-case instance. In spite of seemingly strong restrictions that we impose, the ratio of our algorithm is not far from the best achievable bound by any deterministic algorithm, which we show is at least $\frac{9+\sqrt{17}}{8}\approx 1.640$. Altogether, we obtain a near-complete picture of competitiveness for $m=2$.
The remainder of the paper demonstrates a disparity between increasing the number $m$ of machines and increasing the number $K$ of scenarios. For a fixed but arbitrary $m$, we show that there exists a sufficiently large number $K(m)$ of scenarios and a family of instances with $K(m)$ scenarios, on which no deterministic online algorithm can achieve a ratio better than $m$. Furthermore, we can even choose all instances to have unit processing times $p_{j}\equiv 1$. This result is particularly noteworthy, considering that any online algorithm on $m$ machines is trivially $m$-competitive: Indeed, the online output is at most the total load and an offline optimum is at least the average load, and these are by at most a factor of $m$ apart.
To find the suitable number $K(m)$, we interpret the instances as hypergraphs, passing to a problem we call Online Makespan Hypergraph Coloring$(m)$ instead (cf. Section 1.2). Strikingly, such a family can be created using hypertrees on which no deterministic algorithm can achieve a coloring without a monochromatic edge simultaneously.
@@
Let $m\in\mathbb{N}$. There exists a number $K(m)$ such that for any $r<m$, no deterministic algorithm for Online Makespan Hypergraph Coloring$(m)$ is $r$-competitive, even when restricted to hypertrees with $K(m)$ hyperedges.
The statement of Theorem 1.2 is best possible in the sense that the numbers $K(m)$ cannot be replaced by a global $K$ for which the incompetitiveness holds for all $m$ (cf. Theorem 2.3).
@@
[…]
@@
2 $>1.640$, $\leq\nicefrac{{5}}{{3}}$ $=2$ $=2$
3 $\leq 2$ $\geq 2$, $\geq 2$
$=2$ for $p_{j}\equiv 1$ $=3$ for $K\geq 233$ and $p_{j}\equiv 1$
$\geq 2$, $<4$ $\geq 2$, $<K+1$
4+ $\geq 3$ $=2$ for $p_{j}\equiv 1$ $=m$ for $K$ sufficiently large
and $p_{j}\equiv 1$
@@ 1.2 Preliminaries
Throughout this paper, for $n\in\mathbb{N}$, we denote by $[n]\coloneqq\{1,\ldots,n\}$ the set of numbers $1$ through $n$. An instance $I=(n,(p_{j})_{j\in[n]},\{S_{k}\}_{k\in[K]})$ of Online Makespan Scheduling under Scenarios$(m,K)$ (OMSS$(m,K)$) is parametrized by a number $m\in\mathbb{N}$ of machines and a number $K\in\mathbb{N}$ of scenarios, and defined as a tuple consisting of a number $n\in\mathbb{N}$ of jobs, a processing time $p_{j}\geq 0$ for each job $j$ and job subsets $S_{1},\ldots,S_{K}\subseteq[n]$, called scenarios. As we usually work with a single instance at a time, we often suppress this notation in the sequel as well as the parameters $m$ and $K$, which are often specified per section.
For a set $S\subseteq[n]$, we denote $p(S)\coloneqq\sum_{j\in S}p_{j}$. Our objective is to find a partitioning $J_{1}\dot{\cup}\ldots\dot{\cup}J_{m}$ of the set of jobs that minimizes $\max_{k\in[K]}\max_{i\in[m]}p(J_{i}\cap S_{k}).$ In other words, we consider schedules that arise from restricting a partition to each of the scenarios, and our goal is then to minimize the makespan of the worst-case scenario, which we also call the makespan of the schedule. Furthermore, our instance is online in the sense that parameters $m$ (number of machines) and $K$ (the number of scenarios) are known in advance, but jobs $j\in[n]$ as well as their processing times $p_{j}$ and the indices $k\in[K]$ such that $j\in S_{k}$ are revealed one job $j$ at a time. Once the information of a job $j$ is revealed in addition to already known information about the jobs in $[j-1]$, the job $j$ must be assigned to a machine $i\in[m]$, which is an irrevocable decision.
In our analyses and arguments, we use the notion of a completion time of a job $j$. In classical scheduling, this is defined as the sum of processing times preceding (including) $j$ that are on the same machine as $j$. In our setting, the restricted schedules are evaluated without idle times, meaning that such completion times might be different for different scenarios. Accordingly, we define the completion time of job $j\in J_{i}\cap S_{k}$ in scenario $S_{k}$ as $C^{k}_{j}\coloneqq\sum_{j^{\prime}\in J_{i}\cap S_{k}\cap[j]}p_{j^{\prime}}$ and the completion time as $C_{j}\coloneqq\max_{k\in[K],j\in S_{k}}C_{j}^{k}$.
@@
We also consider the special case of unit processing times $p_{j}\equiv 1$. In this special case, we may view our jobs as nodes, our scenarios as hyperedges containing a subset of the nodes and the partition $J_{1}\dot{\cup}\ldots\dot{\cup}J_{m}$ as a coloring $c\colon[n]\to[m]$ of the nodes. Then, our problem can be formulated as an online hypergraph coloring problem, in which we color a hypergraph $H=(V,E)$ with a fixed number of colors $1,\ldots,m$ and seek to minimize $\max_{i\in[m],e\in E}|\{c^{-1}(i)\}\cap e|$. The nodes $j\in[n]$ are revealed one at a time in increasing order along with partial hyperedges $e\cap[j]$ $(e\in E,j\in[n])$. [^2] According to this definition, hyperedges can have size $1$, although such hyperedges can and will be omitted as they pose no difference with respect to the objective function. We shall name this problem Online Makespan Hypergraph Coloring$(m)$ (OMHC$(m)$) and address it in Section 4.
@@ 1.3 Related Work
A problem similar to Online Makespan Hypergraph Coloring, often called Online Hypergraph Coloring, has been studied extensively over the years. In this problem, the objective is to use the smallest number of colors while avoiding a monochromatic hyperedge, i.e., a hyperedge whose nodes all attain the same color. For hypertrees with $n$ nodes, an $O(\log(n))$-competitive algorithm is known to be best possible ([hypertreecoloring]), while for general graphs with $n$ nodes, no $o(n)$-competitive algorithm can exist ([NAGYGYORGY200823]). However, this standard model differs from ours, as a hyperedge is revealed only once all of its nodes have been revealed. This nuance deems it much easier to construct families of instances to achieve lower bounds. Moreover, due to differences in their respective objective functions, this problem does not seem to have a direct relation to Online Makespan Hypergraph Coloring. However, our approach has implications for a variant of Online Hypergraph Coloring in which partial hyperedges are known (see Section 1.1).
@@ 2 Extending Graham’s List Scheduling
Our first observation is a simple algorithm that generalizes Graham’s List Scheduling Algorithm for OMSS$(m,K)$ and performs well when there are too many machines to force the schedule to create significant discrepancy. Our algorithm beats the trivial bound of $m$ if $m>K+1$.
We first observe that if the number $m$ of machines is large enough, we can apply the pigeonhole principle to find a machine that is not among the most loaded machines of any scenario. For $s\in[m-1]$ and $k\in[K]$, we call a machine $i$ $s$-favorable with respect to the $k$-th scenario if there are at least $s$ distinct machines $i^{\prime}\neq i$ that each have at least the load of machine $i$ in the $k$-th scenario, i.e., $p(J_{i^{\prime}}\cap S_{k})\geq p(J_{i}\cap S_{k})$.
@@
Notice that the aforementioned competitive ratio is strictly smaller than $K+1$.
This algorithm is also applicable in the special case where every job appears in at most $k$ scenarios for a fixed $k\in\mathbb{N}$, yielding a ratio of $\nicefrac{{(m-1)}}{{\left\lceil\frac{m}{k}\right\rceil}}+1$. Furthermore, plugging in $K=1$ yields the well-known competitive ratio of $2-\nicefrac{{1}}{{m}}$ of Graham’s List Scheduling Algorithm.
@@ 3 Near-Tight Bounds for $m=2$ Machines
On the other hand, this statement cannot be extended to $K=2$. To show this, we present a minimalistic yet insightful algorithm (Algorithm 1) for makespan scheduling under $K=2$ scenarios and $m=2$ machines. Our general idea is to fix a simple rule for assigning double-scenario jobs, i.e., jobs in $S_{1}\cap S_{2}$, and partition the single-scenario jobs $j\in S_{1}\triangle S_{2}$ so that the machines are reasonably balanced even after upcoming hypothetical double-scenario jobs.
@@
If $j\in S_{1}\cap S_{2}$, then assign the job $j$ to a machine $i$ such that the makespan of the schedule was previously attained by the other machine $3-i$. In case of a tie, select the machine that received the job $j-1$ ($j=1$ is assigned to the first machine if applicable).
@@
$\rho=\frac{\max_{i,k\in\{1,2\}}p(S_{k}\cap J_{i})}{\max\left\{\max_{k\in\{1,2\}}\left\{\frac{p(S_{k})}{2}\right\},\max_{j\in[n]\cap\mathrm{argmax_{j\in[n]}\{C_{j}\}}}\{p_{j}\}\right\}},$

### v2

@@ Abstract
For varying values of the number of scenarios and machines, we explore the competitiveness of online algorithms. We prove tight and near-tight bounds, several of which are achieved through novel constructions. In particular, we leverage the interplay between the unit processing time case of our problem and the hypergraph coloring problem both ways: We use hypergraph coloring techniques to steer an adversarial family of instances proving lower bounds, which in turn leads to lower bounds for several variants of online hypergraph coloring.
@@ 1 Introduction
Moreover, even special cases of online makespan scheduling under scenarios correspond to interesting questions in their own regard. The special case $p_{j}\equiv 1$ of unit processing times can be related to some cases of discrepancy minimization as well as hypergraph coloring.
In the large array of previous research on combinatorial optimization under scenarios, it was shown for several offline problems that optimizing under a number of scenarios is computationally harder than their underlying single-scenario counterparts. Accordingly, our goal in this work is to gain a better understanding of the intersection of scenario-based models and competitiveness, in particular how competitiveness becomes harder under a growing number of scenarios.
@@ 1.1 Our Contribution
By a simple example, we show that for any number $m\geq 2$ of machines and (at least) $3$ scenarios, no algorithm with a competitive ratio better than $2$ can exist, even for jobs with unit processing times. To underline the competitiveness gap between two and three scenarios, we contrast the aforementioned result with an algorithm for two machines and two scenarios, which is one of our main results.
@@
This algorithm as well as its analysis combine techniques well-known from algorithms for online makespan scheduling with novel ideas that concern the scenario-based properties of a schedule. What sets our analysis apart from previous work is that we alternate between forward and reverse engineering: Assuming that our algorithm obeys some fixed rules, we explore what we may assume about a well-structured worst-case instance. In spite of seemingly strong restrictions that we impose, the ratio of our algorithm is not far from the best achievable bound by any deterministic algorithm, which we show is at least $\frac{9+\sqrt{17}}{8}\approx 1.640$. Altogether, we obtain a near-complete picture of competitiveness for $m=2$.
The remainder of the paper demonstrates a disparity between increasing the number $m$ of machines and increasing the number $K$ of scenarios. For a fixed but arbitrary $m$, we show that there exists a sufficiently large number $K(m)$ of scenarios and a family of instances with $K(m)$ scenarios on which no deterministic online algorithm can achieve a ratio better than $m$. Furthermore, we can even choose all instances to have unit processing times $p_{j}\equiv 1$. This result is particularly noteworthy, considering that any online algorithm on $m$ machines is trivially $m$-competitive: Indeed, the online output is at most the total load and an offline optimum is at least the average load, and these are by at most a factor of $m$ apart.
To find the suitable number $K(m)$, we interpret the instances as hypergraphs, passing to a problem we call Online Makespan Hypergraph Coloring(m)(m) instead (cf. Section 1.2). Strikingly, such a family can be created using hypertrees on which no deterministic algorithm can achieve a coloring without a monochromatic edge simultaneously.
@@
Let $m\in\mathbb{N}$. There exists a number $K(m)$ such that for any $r<m$, no deterministic algorithm for Online Makespan Hypergraph Coloring(m)(m) is $r$-competitive, even when restricted to hyperforests with $K(m)$ hyperedges.
The statement of Theorem 1.2 is best possible in the sense that the numbers $K(m)$ cannot be replaced by a global $K$ for which the incompetitiveness holds for all $m$ (cf. Corollary 2.5).
@@
An advantage of our approach while proving Theorem 1.2 is that it only features hyperforests. More precisely, we implicitly show that for every $m\in\mathbb{N}$, there is a number $n\leq(2m+1)\uparrow\uparrow 5$ of nodes such that no deterministic online algorithm on $n$ nodes that uses $m$ colors can avoid a monochromatic hyperedge. In other words, in order to color a hyperforest with an arbitrary number $n$ of nodes online without monochromatic edges, we need $\Omega(\mathrm{slog_{5}}n)$ colors (cf. Preliminaries for notation), even if the restrictions $e\cap\{1,\ldots,j\}$ of hyperedges $e$ to revealed nodes are known upon the revelation of nodes $\{1,\ldots,j\}$. Here, the emphasis is on the additional knowledge of partial hyperedges: This makes an adversarial revelation of hyperedges more challenging and to the best of our knowledge, none of the previous lower bound constructions [NAGYGYORGY200823][hypertreecoloring] are applicable under such additional information. This result addresses a completely open problem from [NAGYGYORGY200823] and sets a non-constant lower bound for Online Hypergraph Coloring for hypertrees under the knowledge of partial hyperedges.
@@
2 $\in\left(1.640,{{\frac{5}{3}}}\right]$ =2{=2} =2{=2}
3 $\in\left[\frac{5}{3},2\right]$}
$[2,3]$
=2{=2} for $p_{j}\equiv 1$}
$\in[2,3]$ =3=3 even for $p_{j}\equiv 1$, if $K\geq 233$
4+ $\in\left[\frac{5}{3},3\right)$
$\in[2,4)$
=2{=2} for $p_{j}\equiv 1$}
$\in[2,\min\{m,K+1\})$ =m=m even for $p_{j}\equiv 1$, if $K$ large enough
@@ 1.2 Preliminaries
Throughout this paper, for $n\in\mathbb{N}$, we denote by $n]\coloneqq\{1,\ldots,n\}$ the set of numbers $1$ through $n$. An instance $I=(n,(p_{j})_{j\in[n]},\{S_{k}\}_{k\in[K]})$ of Online Makespan Scheduling under Scenarios(m,K)(m,K) (OMSS(m,K)(m,K)) is parametrized by a number $m\in\mathbb{N}$ of machines and a number $K\in\mathbb{N}$ of scenarios, and defined as a tuple consisting of a number $n\in\mathbb{N}$ of jobs, a processing time $p_{j}\geq 0$ for each job $j$ and job subsets $S_{1},\ldots,S_{K}\subseteq[n]$, called scenarios. As we usually work with a single instance at a time, we often suppress this notation in the sequel as well as the parameters $m$ and $K$, which are often specified per section.
For a set $S\subseteq[n]$, we denote $p(S)\coloneqq\sum_{j\in S}p_{j}$. Our objective is to find a partitioning $J_{1}\dot{\cup}\ldots\dot{\cup}J_{m}$ of the set of jobs that minimizes $\max_{k\in[K]}\max_{i\in[m]}p(J_{i}\cap S_{k}).$ In other words, we consider schedules that arise from restricting a partition to each of the scenarios, and our goal is then to minimize the makespan of the worst-case scenario, which we also call the makespan of the schedule. Furthermore, our instance is online in the sense that the number $m$ of machines is known in advance, but jobs $j\in[n]$ as well as their processing times $p_{j}$ and the indices $k\in[K]$ such that $j\in S_{k}$ are revealed one job $j$ at a time. Once the information of a job $j$ is revealed in addition to already known information about the jobs in [j1][j-1], the job $j$ must be assigned to a machine $i\in[m]$, which is an irrevocable decision.
In our analyses and arguments, we use the notion of a completion time of a job $j$. In classical scheduling, this is defined as the sum of processing times preceding (including) $j$ that are on the same machine as $j$. In our setting, the restricted schedules are evaluated without idle times, meaning that such completion times might be different for different scenarios. Accordingly, we define the completion time of job $j\in J_{i}\cap S_{k}$ in scenario $S_{k}$ as $C^{k}_{j}\coloneqq\sum_{j^{\prime}\in J_{i}\cap S_{k}\cap[j]}p_{j^{\prime}}$ and the completion time as $C_{j}\coloneqq\max_{k\in[K],j\in S_{k}}C_{j}^{k}$. We can further define
$\mathrm{makespan}(j)\coloneqq\max_{j^{\prime}\in[j]}C_{j^{\prime}}.$
@@
We also consider the special case of unit processing times $p_{j}\equiv 1$. In this special case, we may view our jobs as nodes, our scenarios as hyperedges containing a subset of the nodes and the partition $J_{1}\dot{\cup}\ldots\dot{\cup}J_{m}$ as a coloring $c\colon[n]\to[m]$ of the nodes. Then, our problem can be formulated as an online hypergraph coloring problem, in which we color a hypergraph $\mathcal{H}=(V,E)$ with a fixed number of colors $1,\ldots,m$ and seek to minimize $\max_{i\in[m],e\in E}|\{c^{-1}(i)\}\cap e|$. The nodes $j\in[n]$ are revealed one at a time in increasing order along with partial hyperedges $e\cap[j]$ $(e\in E,j\in[n])$. [^1] According to this definition, hyperedges can have size $1$, although such hyperedges can and will be omitted as they pose no difference with respect to the objective function. We shall name this problem Online Makespan Hypergraph Coloring(m)(m) and address it in Section 4. Our construction in Section 5 is even restricted to hyperforests. A hyperforest is a hypergraph without a hypercycle, where a hypercycle is defined as a sequence $v_{1},e_{1},v_{2}\ldots,e_{\ell},v_{\ell+1}=v_{1}$ with pairwise distinct hyperedges $e_{t}$ such that $v_{t},v_{t+1}\in e_{t}$ for all $t\in[\ell]$.
Finally, it is worth mentioning some big numbers that we encounter in Section 5. In particular, we use Knuth’s arrow notation to denote $F(x)\coloneqq x\uparrow\uparrow 5\coloneqq x^{x^{x^{x^{x}}}}$ and the superlogarithm $\mathrm{slog}_{5}(x)\coloneqq F^{-1}(x)$.
@@ 1.3 Related Work
A problem similar to Online Makespan Hypergraph Coloring, often called Online Hypergraph Coloring, has been studied extensively over the years. In this problem, the objective is to use the smallest number of colors while avoiding a monochromatic hyperedge, i.e., a hyperedge whose nodes all attain the same color. For hypertrees with $n$ nodes, an $O(\log(n))$-competitive algorithm is known to be best possible ([hypertreecoloring]), while for general graphs with $n$ nodes, no $o(n)$-competitive algorithm can exist ([NAGYGYORGY200823]). However, this standard model differs from ours, as a hyperedge is revealed only once all of its nodes have been revealed. This nuance deems it much easier to construct families of instances to achieve lower bounds. Moreover, due to differences in their respective objective functions, this problem does not seem to have a direct relation to Online Makespan Hypergraph Coloring. However, our approach conveniently enables implications for a variant of Online Hypergraph Coloring in which partial hyperedges are known (see Section 1.1).
@@ 2 Extending Graham’s List Scheduling
Our first observation is a simple algorithm that generalizes Graham’s List Scheduling Algorithm for OMSS(m,K)(m,K) and performs well when there are too many machines to force the schedule to create significant disbalance. Our algorithm beats the trivial bound of $m$ if $m>K+1$.
We first observe that if the number $m$ of machines is large enough, we can apply the pigeonhole principle to find a machine that is not among the most loaded machines of any scenario. For $s\in[m-1]$ and $k\in[K]$, we call a machine $i$ $s$-favorable with respect to the $k$-th scenario if there are at least $s$ machines $i^{\prime}\neq i$ that have at least the load of machine $i$ in the $k$-th scenario, i.e., $p(J_{i^{\prime}}\cap S_{k})\geq p(J_{i}\cap S_{k})$.
@@
Proof 2.4.
Let $i\in[m]$ be a machine and $k\in[K]$ a scenario that attains the makespan of the schedule, and let $j\in[n]$ be the smallest index such that $p(J_{i}\cap S_{k}\cap[j])=\mathrm{makespan}(n)$. We define $L\coloneqq p(J_{i}\cap S_{k}\cap[j-1])=\mathrm{makespan}(n)-p_{j}$. By the choice of $i$ upon the insertion of $j$, there are at least $\left\lceil\frac{m}{K}\right\rceil$ machines $i^{\prime}$ with $p(J_{i^{\prime}}\cap S_{k}\cap[j-1])\geq L$, including $i$, so that the offline optimum is bounded from below by
$\operatorname{\mathrm{offl}}$ $\geq\max\left\{p_{j},\frac{p(S_{k})}{m}\right\}\geq\max\left\{p_{j},\frac{p(S_{k}\cap[j-1])+p_{j}}{m}\right\}$ $\geq\max\left\{p_{j},\frac{\left\lceil\frac{m}{K}\right\rceil\cdot L+p_{j}}{m}\right\}.$
Hence, for $p_{j}\geq\frac{\left\lceil\frac{m}{K}\right\rceil}{m-1}L$, the competitive ratio $\rho$ is bounded by
$\rho\leq\frac{L+p_{j}}{p_{j}}\leq 1+\frac{m-1}{\left\lceil\frac{m}{K}\right\rceil},$
and otherwise it is also bounded by
$\rho\leq\frac{L+p_{j}}{\frac{\left\lceil\frac{m}{K}\right\rceil\cdot L+p_{j}}{m}}=m\cdot\frac{L+p_{j}}{\left\lceil\frac{m}{K}\right\rceil\cdot L+p_{j}}\leq m\cdot\frac{1+\frac{\left\lceil\frac{m}{K}\right\rceil}{m-1}}{\left\lceil\frac{m}{K}\right\rceil+\frac{\left\lceil\frac{m}{K}\right\rceil}{m-1}}=1+\frac{m-1}{\left\lceil\frac{m}{K}\right\rceil}.$
Here, the second inequality holds since $\left\lceil\frac{m}{K}\right\rceil>1$.
Corollary 2.5.
The above algorithm is $(K+1)$-competitive.
This algorithm is also applicable in the special case where every job appears in at most $k$ scenarios for a fixed $k\in\mathbb{N}$, yielding a ratio of $\frac{m-1}{\left\lceil\frac{m}{k}\right\rceil}+1$. Furthermore, plugging in $K=1$ yields the well-known competitive ratio of $2-\frac{1}{m}$ of Graham’s List Scheduling Algorithm.
@@ 3 Near-Tight Bounds for $m=2$ Machines
On the other hand, this statement cannot be extended to $K=2$. To show this, we present a minimalistic yet insightful algorithm (Algorithm 1) for makespan scheduling under $K=2$ scenarios and $m=2$ machines. To this end, we combine ideas about what the algorithm can feature with observations about how a worst-case instance is bound to look under the ideas we have fixed. To this end, we first need to introduce a few notions and machinery.
3.1 Preparation for the Algorithm
Fixing a rule.
Our general idea is to fix a simple rule for assigning double-scenario jobs, i.e., jobs in $S_{1}\cap S_{2}$, and partition the single-scenario jobs $j\in S_{1}\triangle S_{2}$ so that the machines are reasonably balanced even after upcoming hypothetical double-scenario jobs.
@@ 3.1 Preparation for the Algorithm
If $j\in S_{1}\cap S_{2}$, then assign the job $j$ to a machine $i$ such that the makespan of the schedule was previously attained by the other machine 3i3-i. In case of a tie, select the machine that received the job j1j-1 ($j=1$ is assigned to the first machine if applicable).
A more approachable notion of competitiveness.
@@
$\rho=\frac{\max_{i,k\in\{1,2\}}p(S_{k}\cap J_{i})}{\max\left\{\max_{k\in\{1,2\}}\left\{\frac{p(S_{k})}{2}\right\},\max_{j\in[n]\cap\mathrm{argmax\{C_{j}\}}}\{p_{j}\}\right\}},$